# Continuous Speech Synthesis using per-token Latent Diffusion

### v1

@@ 3.2 SALAD: Speech synthesis using Autoregressive LAtent Diffusion
SALAD performs zero-shot text to speech, which can synthesize speech based on a given text and a speaker prompt. It does so by predicting the continuous VAE latents $\bm{x}$ with the per-token diffusion head. Our approach utilizes semantic tokens $\bm{w}$ as an auxiliary representation that provides contextual information and determines the stopping condition. We provide two variants for SALAD:
  • Semantic to Acoustic (S2A) - predicts acoustic features based on semantic tokens, and relies on an external text-to-semantic model to produce the semantic tokens (Figure 3).
  • Text to Acoustic (T2A) - predicts acoustic features and semantic features directly from text, relying on the stopping condition of the semantic tokens (Figure 4).
3.2.1 Semantic to Acoustic (S2A)
@@
Following Kharitonov et al. (2023), we divide synthesis into two tasks: text-to-semantic (T2S) and semantic-to-acoustic (S2A), each tackled by a different model. The T2S model predicts the discrete semantic tokens based on the text and speaker autoregressively using a causal transformer:
$p(w_{1},..,w_{n}|t,s)=\prod^{n}_{i=1}p(w_{i}|t,s,w_{1},..,w_{i-1})$
Importantly, the T2S model determines the synthesized audio length. Then, our S2A model predicts the acoustic tokens based on the semantic tokens, using an additional transformer model with the diffusion head to model the continuous distributions. We provide two variants of semantic-to-acoustic models, AR and NAR, following the literature on discrete acoustic token modeling.
Autoregressive S2A
In our AR semantic-to-acoustic model, our training objective is $p(x_{1},...,x_{n}|\bm{w},\bm{s})=\prod^{n}_{i=1}p(x_{i}|\bm{w},\bm{s},x_{1},..,x_{i-1})$. We input the latent frames, the semantic tokens, and speaker prompt into a causal transformer and obtain the contextual condition vectors $z_{1},..,z_{n}=\Theta(\bm{w},\bm{s},x_{1},...,x_{n})$. The frame $x_{i}$ is predicted given $z_{i-1}$ and our loss is:
$\mathcal{L}(\bm{x},\bm{z})=\sum^{n}_{i=1}\mathbb{E}_{\epsilon_{i},t_{i}}\left[\|\epsilon_{i}-\epsilon_{\theta}(x_{i}^{t},t_{i},z_{i-1})\|^{2}\right]$
During inference, the T2S model generates semantic tokens $\hat{\bm{w}}$. Then, the S2A model generates the continuous latent vectors based on the predicted semantic tokens, by computing the contextual embedding $z_{i}$, and uses diffusion head inference to sample the next continuous frame $x_{i+1}$.
Figure 3: Synthesis using Semantic-to-Acoustic models
Figure 4: Synthesis using Text-to-Acoustic models
Non-Autoregressive S2A
MaskGIT (Chang et al., 2022) trains a bidirectional transformer on a discrete masked language modeling objective. During inference, it unmasks tokens over $K$ inference steps, where each step is based on the previously predicted tokens. Soundstorm extended this procedure to predict the RVQ codes based on semantic tokens, by applying the MaskGIT procedure for each RVQ layer. We extend the MaskGIT procedure to predict continuous acoustic tokens based on semantic tokens, using the diffusion head defined in Section 3.1. Given a schedule function $\gamma(r):[0,1]\to[0,1]$ and a sequence $x_{1},..,x_{n}$, we sample $r\in U[0,1]$ and mask out $\gamma(r)\cdot n$ acoustic tokens. Define the random masking indicators $\bm{m}=(m_{1},...,m_{n})$ where $m_{i}\in\{0,1\}$, we replace masked acoustic tokens with a fixed learnable embedding $q$ and define $r_{i}=m_{i}\cdot q+(1-m_{i})\cdot x_{i}$. As in SoundStorm, a semantic token $w_{i}$ and masked acoustic tokens $r_{i}$ are embedded into a single transformer token, and speaker prompt tokens are left unmasked and appended as context. The resulting sequence is fed into a bidirectional transformer, and the resulting contextual vectors $z_{1},..,z_{n}$ are used to predict the masked acoustic tokens using the diffusion head. The loss function is the denoising of the masked acoustic tokens
$\mathcal{L}(\bm{x},\bm{z},\bm{m})=\sum^{n}_{i=1}\mathbb{E}_{\epsilon,t}\left[m_{i}\cdot\|\epsilon-\epsilon_{\theta}(x_{i}^{t},t,z_{i})\|^{2}\right]$
During inference, MaskGIT chooses the tokens with the highest confidence. In a continuous diffusion model, measuring the confidence is not trivial, so we opt to select the tokens to unmask at random. We first check the influence of confidence-based selection in MaskGIT inference in Section A.2, and conclude that random unmasking is surprisingly superior to the confidence-based unmasking, which validates our design choice. Specifically, given a sequence of $n$ tokens, $K$ maskGIT steps, in every step $k\in[K]$ we unmask $n\left(\gamma\left(\frac{k}{K}\right)-\gamma\left(\frac{k-1}{K}\right)\right)$ additional tokens selected at random, by applying the per-token diffusion head.
3.2.2 Text-to-Acoustic (T2A)
@@
Decoupling TTS into T2S and S2A requires training two models of similar size, and applying two steps of inference, which greatly increases the compute requirements and latency. Therefore, we suggest an end-to-end text-to-acoustic model (T2A) which predicts the acoustic features directly from the text and the speaker prompt. The T2A model predicts the semantic and acoustic features in parallel, where the semantic token prediction is an auxiliary task that allows conditioning on contextual information and provides a stopping condition. We add an additional prediction MLP to predict the discrete semantic tokens. We adopt the delay pattern suggested by (Copet et al., 2024) such that every acoustic token $x_{i}$ can be predicted based on the semantic token $w_{i}$. Define $r_{i}=(w_{i},x_{i-1})$, we extract contextual features from our transformer backbone, based on the text and speaker prompt $z_{i}=\Theta(t,s,r_{1},...r_{i})$, which is used to predict $w_{i+1}$ using the cross-entropy loss $L_{s}$ and $x_{i}$ using the diffusion loss $L_{a}$. We weigh the two losses to $\mathcal{L}=\alpha\mathcal{L}_{a}+(1-\alpha)\mathcal{L}_{s}$. We halt the generation after the semantic prediction head samples an EOS token. We note that the audio duration is predicted on the-fly based on the model’s predictions, unlike most diffusion-based TTS models, where the audio duration is predetermined.
3.3 Discrete Baselines
All SALAD models use common discrete architectures and only replace the input projection and prediction head, so we can implement a discrete variant for each method proposed in Section 3.2. The discrete methods use an RVQ-GAN quantizer, which yields a sequence of discrete codes $(q^{1},...,q^{Q})$ for each frame. These codes are predicted from the contextual embedding $z$ by MLP prediction heads, one for each codebook. The S2A-AR discrete model predicts all codes in parallel, using the delay pattern proposed in MusicGen (Copet et al., 2024). The S2A-NAR discrete model implements SoundStorm (Borsos et al., 2023a), which applies $Q$ MaskGIT procedures, one for each codebook. Unlike the vanilla SoundStorm implementation, we replace the confidence-based unmasking with random unmasking, as we have found it to be superior. Given $K$ MaskGIT steps, SoundStorm requires $QK$ passes through the transformer, as it employs a MaskGIT procedure per RVQ layer, unlike SALAD-NAR, which requires $K$ transformer passes. The T2A discrete model predicts semantic and acoustic tokens in parallel from text, using the delay pattern described above. We use the MusicGen parallel prediction method, treating the semantic tokens as the coarsest codes. In our listening test, we compare to the external XTTS (Casanova et al., 2024), a commercial SOTA zero-shot TTS model.
4 Experiments
4.1 Experimental Setup
Datasets
@@ 4.1 Experimental Setup
We train all models on the English subset of multi-lingual LibriSpeech (MLS) (Pratap et al., 2020), which contains 10M examples of 10-20 seconds, resulting in 45K hours. To avoid over-exposure of a few speakers, we limit the maximal number of utterances per speaker to 10K, resulting in 5.2M examples. We evaluate all models on LibriSpeech test-clean (Panayotov et al., 2015), which consists of 2620 utterances by 40 speakers. All speakers in the test set are excluded from the training set. We filter the dataset to utterances with lengths of 8-25 seconds, and then limit to at most 15 samples per speaker, resulting in 564 utterances for evaluation.
Tokenization
@@
To derive acoustic tokens, we train continuous $\beta$-VAE-GAN, with a varying bottleneck dimension $d\in\left\{8,16,24,32\right\}$, and set the KL-divergence regularization to $\beta=5\cdot 10^{-5}$, as done in Tschannen et al. (2023). We also train discrete RVQ-GAN models with $q\in\left\{4,8,12\right\}$ codebooks, each with 1024 entries. In addition, we apply quantizer dropout (Zeghidour et al., 2021) with $p=0.5$. All compression models are trained on MLS-English, DAPS, LibriTTS, LibriTTS-R and LJ-Speech, which balance between high and mid quality recordings (Shechtman & Dekel, 2024). The all-training hyperparameters follow the original recipe proposed by Kumar et al. (2024). We extract semantic tokens by quantizing the embeddings of the 11th layer of W2V-BERT (Barrault et al., 2023) using minibatch K-means with 1024 centroids. We further compress the semantic tokens using a BPE tokenizer with a vocabulary of 16384 tokens. This is done to shorten the sequence and balance the tokens’ distribution (Dekel & Fernandez, 2024). We note that only the T2S model leverages the BPE-compressed semantic tokens, as well as the the S2A-AR models, as other models embed semantic and acoustic features into a shared token space. We also train a text BPE tokenizer on the transcripts of our training set, with a vocabulary of 16384 tokens.
Architecture
@@
We use a transformer backbone with $d=1024$, $d_{ff}=4096$, $24$ layers, $16$ heads, sinusoidal positional embedding, GeLU activation, and a dropout rate of $0.1$, resulting in models with roughly 350M parameters. VAE embeddings are projected using a linear layer, while RVQ tokens are embedded using $Q$ lookup tables, which are summed into a single embedding. We use Classifier-Free Guidance (Ho & Salimans, 2022) and randomly omit the speaker prompt with $p=0.1$ during training. In the MaskGIT NAR experiments, we use the cosine masking schedule, and apply a total of $64$ inference steps, where the SoundStorm model with $4$ codebooks performs $16$ inference steps per layer. RVQ codes are predicted using a $Q$ MLP heads with four hidden layers. We sample from discrete distributions using Top $k=10$ sampling, with a temperature of $\tau=1$, a repetition penalty of $1.05$, and a CFG scale of $\alpha=3$.
@@
We use a diffusion process with $T=1000$ steps, where betas are logarithmicly spaced between $\beta_{0}=2e-4$ and $\beta_{T}=0.03$. Our per-token diffusion head is an MLP network with 12 residual layers, that predicts the noise $\epsilon$ given the transformer embedding vector $z$, the noisy input $x^{t}$, and the diffusion step $t$. Each residual block consists of layer normalization, linear layer, SiLU activation, and dropout with $p=0.1$. During inference, we apply $20$ diffusion steps for sampling, with a default noise scale of $1$. We use the AdamW optimizer, with $lr_{max}=3e-4$ and $lr_{min}=3e-5$, weight decay $0.1$, and a clip gradient norm of $1$, and train with FP16 mixed precision. We linearly warm up the learning rate from $lr_{min}$ across 32K iterations to $lr_{max}$ and decay the learning rate back to $lr_{min}$ over 300K steps using a cosine schedule. Each global batch size has approximately 150K acoustic tokens (200 samples). Each model was trained with 8 A100 80GB GPUs.
Metrics
We measure Audio Quality using UTMOS (Saeki et al., 2022) which produces a quality score in the range of 1-5 (higher is better). Intelligibility is measured by the character error rate (CER) in percentages (%) between the ground-truth text and the Whisper transcripts (Radford et al., 2023) of the synthesized audio. Speaker Similarity is measured by the cosine similarity to the prompt, comparing the embedding of WavLM-TDNN (Chen et al., 2022), a popular speaker verification model. This metric was also reported in Vall-E and subsequent studies (Wang et al., 2023; Chen et al., 2024). The similarity score predicted is in the range of $[-1,1]$, where a larger value indicates a higher similarity.
For the subjective Listening Tests, we selected one random utterance for every speaker in LibriSpeech test-clean (20 female and 20 male speakers), resulting in 40 utterances for evaluation. The selected utterances were confined to have at most 200 characters to enable the comparison with XTTS xttsv2 (Casanova et al., 2024) (xttsv2 demo limitation). For each sample, we selected a three-second-long speaker prompt from another random utterance of the same speaker. Each system synthesizes the desired utterance based on the same text and speaker prompt. All experiments were conducted on the Amazon Mechanical Turk (AMT) crowd-sourcing platform with votes collected from 39-58 subjects qualified as masters (Sodré & Brasileiro, 2017).
In the first Listening Test we assess speech quality and naturalness by the standard 5-point scale Mean Opinion Score (MOS)  (Ribeiro et al., 2011). 25 distinct subjects assessed each utterance. We report the average scores and the 95% confidence interval. In the second Listening Test we asses the Speaker Similarity by a 4-level pairwise similarity test, as in (Wester et al., 2016; Kons et al., 2018), where subjects were presented with (utterance, prompt) pairs and asked to rank speaker similarity of each pair on a 4-level categorical scale (definitely different speakers, probably different speakers, probably the same speaker, definitely the same speaker). Each utterance was assessed by 20 distinct subjects on average. We report the mean similarity score and the 95% confidence interval while attaching 1-4 numerical values to the above categories, as in (Kons et al., 2018).
Task Modeling Representation UTMOS $\uparrow$ STT CER (%) $\downarrow$ Similarity $\uparrow$
Ground Truth 4.121 0.528 0.736
Text to Acoustic AR Continuous 4.280 0.739 0.539
Text to Acoustic AR Discrete 4.270 2.298 0.600
Semantic to Acoustic AR Continuous 4.27 2.198 0.588
Semantic to Acoustic AR Discrete 4.348 1.231 0.549
Semantic to Acoustic NAR Continuous 4.277 1.393 0.558
Semantic to Acoustic NAR Discrete 4.351 1.846 0.602
Table 1: Objective evaluation of LibriSpeech test-clean
@@
4.2 Results
Objective Evaluation
We evaluate all models on zero-shot TTS. Given a text and a three-second speaker prompt, which is taken randomly from another utterance of the same speaker, the model attempts to synthesize the audio with the identity and prosody similar to the prompt. All models use the same random prompt for each sample. We compare two variants of models that perform Semantic to Acoustic (S2A), and another variant that performs Text to Acoustic (T2A) directly. When using S2A models, we first run the Text to Semantic (T2S) model and use the predicted semantic tokens, as in Figure 3. The discrete models rely on a $4$ codebook model, while the continuous make use of a $d=8$ VAE embedding. Table 1 shows that the continuous models are competitive with their discrete benchmarks. The continuous T2A model presents the highest intelligibility score, making it the most reliable model when having to synthesize an exact text. However, the speaker similarity scores of the discrete T2A and S2A-NAR model are higher. We note that in cases of accented speech or low quality recordings, when the speaker similarity increases, the intelligibility and audio quality often decreases. We did not report objective scores for XTTS due to the sample limit in their demo.
Subjective Evaluation
We conduct the two subjective listening tests, described above, to compare the following systems: (1) Ground Truth audio (2) XTTSv2 (Casanova et al., 2024) (3) T2A Continuous (4) T2A Discrete (5) S2A NAR Continuous (6) S2A NAR Discrete. Figure 5(a) reports the mean opinion score (MOS) results, suggesting that the difference between the ground-truth audio (GT) to both T2A continuous model and the NAR discrete model is statistically insignificant ($p>0.01$). Figure 5(b) presents the speaker similarity average score with 95% confidence intervals, suggesting similar or better speaker similarity scores for all the systems but XTTSv2. More precise analysis with two-sided Wilkinson rank-sum test (Wilcoxon, 1945) reveals that both the T2A continuous and the NAR discrete models do not differ ($p>>0.01$) from the GT in terms of speaker similarity, while the T2A discrete model is marginally better than the GT ($p=0.0105$). The NAR continuous model, however, is marginally worse than the GT ($p=0.0111$).
(a) CFG scale
(b) Noise scale
(c) Diffusion steps
(d) MaskGIT steps
Figure 6: Inference hyperparameters influence
4.3 Ablation Study
Inference Hyperparameters
We now turn to investigate the influence of inference hyperparameters on synthesized speech. We used the T2A model to investigate classifier-free guidance (CFG), noise scale and the number diffusion steps, and the S2A-NAR model for the MaskGIT inference experiment. In every experiment, we fix all values to the default inference values following those described in Section 4, and change only a single hyperparameter. The CFG linear extrapolation coefficient increases the speaker similarity, but degrades the automatic quality metric, as seen in Figure 6(a). Next, we scale the noise level added in each diffusion step by scaling the $\beta_{t}\epsilon$ term in Equation 2, and see improvements in similarity but degradation in the UTMOS quality score (Figure 6(b)). We also examine the number of diffusion steps, which improve similarity until reaching 20 diffusion steps, and also degrade UTMOS (Figure 6(c)). The number of MaskGIT in the NAR model shows consistent improvement in both the speaker similarity and UTMOS (Figure 6(d)).
High-Fidelity Modeling
When increasing the number of RVQ codebooks or the VAE embedding dimension, the reconstruction quality increases, but language modeling can be difficult (Shen et al., 2023). Figure 7(a) shows the reconstruction quality measured by PESQ (Rix et al., 2001), for different numbers of RVQ codebooks and VAE embedding dimensions. One concern regarding RVQ modeling is that the fine codes quantize noise, leading to a high gradient contribution of random classification problems. We measure the noise sensitivity per codebook by adding Gaussian noise into raw samples, compressing them with the RVQ model, and checking the ratio of change per codebook. Results in Figure 7(b) suggest that fine codebooks are indeed more sensitive to noise. In Figure 7(c), we calculate per-codebook validation cross-entropy loss in the discrete 12-codebook T2A-AR model, suggesting the model struggles to reduce uncertainty in finer codebooks. This phenomenon occurs despite the delay-pattern described in Section 3, where the finer RVQ levels are conditioned on coarser layers of the same frame. Finally, we compare the generation quality with less-compressed representations. The results in Table 2 show that when increasing the fidelity, the intelligibility drop of the high-fidelity discrete T2A is greater comparing to the continuous model.

### v2

@@ IV SALAD
SALAD is an end-to-end zero-shot TTS model, that given text and a speaker prompt autoregressively predicting continuous VAE latents $\boldsymbol{x}$ using per-token latent diffusion. Our method utilizes semantic tokens $\boldsymbol{w}$ as an auxiliary representation that provides contextual information and determines the stopping condition. We project both the semantic and acoustic representations into a shared space, and sum them to obtain a fused representation. We adopt the delay pattern suggested by [Musicgen], such that each acoustic token $x_{i}$ is predicted based on the semantic token $w_{i}$, for which we define $r_{i}=f(w_{i},x_{i-1})$.
@@
Formally, model predicts both discrete semantic tokens and continuous acoustic tokens autoregressively, based on the text and speaker, using a causal transformer:
$p\!\big((w_{1},x_{0}),\ldots,(w_{n},x_{n-1})\mid t,s\big)\\ =\prod_{i=1}^{n}p\!\big((w_{i},x_{i-1})\mid t,s,r_{1},\ldots,r_{i-1}\big).$ (3)
@@
For each autoregressive step $i$, we extract contextual features from our transformer backbone, based on the text and speaker prompt $z_{i}=\Theta(t,s,r_{1},...r_{i})$, which is used to predict $w_{i+1}$ using the cross-entropy loss $L_{s}$ by the MLP head, and $x_{i}$ using the diffusion loss $L_{a}$ by the diffusion head (see the system overview in Figure 3). We halt the generation once the semantic prediction head samples an EOS token. We note that audio duration is predicted on the-fly based on the model’s predictions, unlike most diffusion-based TTS models, where the audio duration is predetermined.
V Experiments
V-A Experimental Setup
@@
Datasets.  We train our models on the English subset of multi-lingual LibriSpeech (MLS) [pratap2020mls], which contains 10M examples of 10-20 seconds, resulting in 45K hours. To avoid over-exposure of a few speakers, we limit the maximal number of utterances per speaker to 10K, resulting in 5.2M examples. We evaluate all models on LibriSpeech test-clean [panayotov2015librispeech], which consists of 2620 utterances by 40 speakers. All speakers in the test set are excluded from the training set. We filter the dataset to utterances with lengths of 8-25 seconds, and then limit to at most 15 samples per speaker, resulting in 564 utterances for evaluation.
@@
Tokenization.  To derive acoustic tokens, we train continuous $\beta$-VAE-GAN, with a varying bottleneck dimension $d\in\left\{8,16,24,32\right\}$, and set the KL-divergence regularization to $\beta=5\cdot 10^{-5}$, as done in [GIVT]. We also train discrete RVQ-GAN models with $q\in\left\{4,8,12\right\}$ codebooks, each with 1024 entries. In addition, we apply quantizer dropout [Soundstream] with $p=0.5$. All compression models are trained on MLS-English, DAPS, LibriTTS, LibriTTS-R and LJ-Speech, which balance between high and mid quality recordings [shechtman24_interspeech]. The all-training hyperparameters follow the original recipe proposed by [DAC]. We extract semantic tokens by quantizing the embeddings of the 11th layer of W2V-BERT [barrault2023seamless] using minibatch K-means with 1024 centroids.
@@
Architecture.  We use a transformer backbone with $d=1024$, $d_{ff}=4096$, $24$ layers, $16$ heads, sinusoidal positional embedding, GeLU activation, and a dropout rate of $0.1$, resulting in models with roughly 350350M parameters. VAE embeddings are projected using a linear layer, while RVQ tokens are embedded using $Q$ lookup tables, which are summed into a single embedding. We use Classifier-Free Guidance (CFG) [CFG] and randomly omit the speaker prompt with $p=0.1$ during training.
RVQ codes are predicted using a $Q$ MLP heads with four hidden layers. For discrete distributions we apply top $k=10$ sampling, with a temperature of $\tau=1$, a repetition penalty of $1.05$, and a CFG scale of $\alpha=3$.
@@
We use a diffusion process with $T=1000$ steps, where betas are logarithmicly spaced between $\beta_{0}=2e-4$ and $\beta_{T}=0.03$. Our per-token diffusion head is an MLP network with 12 residual layers, that predicts the noise $\epsilon$ given the transformer embedding vector $z$, the noisy input $x^{t}$, and the diffusion step $t$. Each residual block consists of layer normalization, linear layer, SiLU activation, and dropout with $p=0.1$. During inference, we apply $20$ diffusion steps for sampling, with a default noise scale of $1$. We use the AdamW optimizer, with $lr_{max}=3e-4$ and $lr_{min}=3e-5$, weight decay $0.1$, and a clip gradient norm of $1$, and train with FP16 mixed precision. We linearly warm up the learning rate from $lr_{min}$ across 32K iterations to $lr_{max}$ and decay the learning rate back to $lr_{min}$ over 300K steps using a cosine schedule. Each global batch size has $\sim$150K acoustic tokens. Each model was trained with 8 A100 80GB GPUs.
@@
Model Type UTMOS CER Sim.
GT 4.121 0.528 0.736
XTTS Disc. 3.91 0.787 0.544
VoiceCraft Disc. 3.986 4.067 0.598
SALAD Disc. 4.27 2.298 0.600
SALAD Cont. 4.28 0.739 0.539
TABLE I: Objective evaluation of LibriSpeech test-clean
Metrics.  We measure Audio Quality using UTMOS [UTMOS] which produces a quality score in the range of 1-5 (higher is better). Intelligibility is measured by the character error rate (CER) in percentages (%) between the ground-truth text and the Whisper transcripts [Whisper] of the synthesized audio. Speaker Similarity is measured by the cosine similarity to the prompt, comparing the embedding of WavLM-TDNN [wavLM], a popular speaker verification model. This metric was also reported in Vall-E and subsequent studies [VALLE, VallE2]. The similarity score predicted is in the range of [1,1][-1,1], where a larger value indicates a higher similarity.
For the subjective listening tests, we selected one random utterance for every speaker in LibriSpeech test-clean (20 female and 20 male speakers), resulting in 40 utterances for evaluation. For each sample, we selected a three-second-long speaker prompt from another random utterance of the same speaker. Each system synthesizes the desired utterance based on the same text and speaker prompt. All experiments were conducted on the Amazon Mechanical Turk crowd-sourcing platform with votes collected from 39-58 subjects qualified as masters [sodre2017analysis].
In the first listening test we assess speech quality and naturalness by the standard 5-point scale Mean Opinion Score (MOS)  [ribeiro2011crowdmos]. 25 distinct subjects assessed each utterance. We report the average scores and the 95% confidence interval.
In the second listening test we asses the Speaker Similarity by a 4-level pairwise similarity test, as in [wester16_interspeech, kons2018neural], where subjects were presented with (utterance, prompt) pairs and asked to rank speaker similarity of each pair on a 4-level categorical scale (definitely different speakers, probably different speakers, probably the same speaker, definitely the same speaker). Each utterance was assessed by 20 distinct subjects on average. We report the mean similarity score and the 95% confidence interval while attaching 1-4 numerical values to the above categories, as in [kons2018neural].
We further compare SALAD with two leading publicly available zero-shot TTS systems: XTTS [XTTS], run with its official package using default inference settings; and VoiceCraft [peng2024voicecraft], using the $330M_{T}TSEnhanced$ model configured for top-k sampling (k = 40), nucleus sampling (p = 1), temperature = 1, and repetition penalty = 3. The VoiceCraft authors report that this configuration outperforms their reimplementation of VALL-E2[VallE2], which we do not include directly in our evaluation, as it is not publicly available and its demo set is too limited for reliable batch metrics.
V-B Results
Subjective Evaluation.  We conduct the two subjective listening tests, described above, to compare the following systems: (1) Ground Truth audio (2) XTTSv2 [XTTS] (3) SALAD (4) SALAD - Discrete. Figure 4(a) reports the mean opinion score (MOS) results, suggesting that the difference between the ground-truth audio (GT) to SALAD is statistically insignificant ($p>0.01$). Figure 4(b) presents the speaker similarity average score with 95% confidence intervals, suggesting similar or better speaker similarity scores for all the systems but XTTSv2. More precise analysis with two-sided Wilkinson rank-sum test [wilcoxon1945individual] reveals that SALAD  do not differ ($p>>0.01$) from the GT in terms of speaker similarity, while SALAD-discrete is marginally better than the GT ($p=0.0105$).
Objective Evaluation.  We evaluate all models on zero-shot TTS. Given a text and a three-second speaker prompt, which is taken randomly from another utterance of the same speaker, the model attempts to synthesize the audio with the identity and prosody similar to the prompt. All models use the same random prompt for each sample.
Table I shows that SALAD achieves superior intelligibility (lowest CER) making it the most reliable model when having to synthesize an exact text, with comparable UTMOS quality to the discrete baseline. However, in terms of speaker similarity SALAD  is comparable to XTTS but the discrete baseline is superior.
(a) CFG scale
(b) Noise scale
(c) Diffusion steps
Figure 6: Inference hyperparameters influence
V-C Ablation Study
High-Fidelity Modeling.  When increasing the number of RVQ codebooks or the VAE embedding dimension, the reconstruction quality increases, but language modeling can be difficult [NaturalSpeech2]. Figure 5(a) shows the reconstruction quality measured by PESQ [PESQ], for different numbers of RVQ codebooks and VAE embedding dimensions. One concern regarding RVQ modeling is that the fine codes quantize noise, leading to a high gradient contribution of random classification problems. We measure the noise sensitivity per codebook by adding Gaussian noise into raw samples, compressing them with the RVQ model, and checking the ratio of change per codebook. Figure 5(b) suggests that fine codebooks are more noise-sensitive. Figure 5(c) shows per-codebook validation cross-entropy loss in the discrete 12-codebook model, indicating difficulty in reducing uncertainty in finer codebooks. This occurs despite the delay pattern in Section IV, where finer RVQ levels are conditioned on coarser layers of the same frame. Finally, we compare the generation quality with less-compressed representations. Table II shows that higher fidelity causes a greater intelligibility drop in the high-fidelity discrete model than in the continuous model.
VAE Sampling  Unlike discrete codebooks or Mel spectrograms, VAE models enable diverse sampling, which may enhance robustness and mitigate the mismatch between training and inference(where prediction rely on previous noisy outputs). To validate this, we compare two SALAD models - one sampling from the VAE distribution $x=\mu+\epsilon\cdot\sigma$ and the other using only the mean $x=\mu$. The results presented in Table III, show a large gap in UTMOS and intelligibility indicating that sampling improves synthesized quality. Listening to VAE-NoSample outputs we observed increasing speaker-inconsistency artifacts, likely due to the absence of sampling noise during training. We suspect this noise improves robustness, and we hypothesize that the high similarity score of VAE-NoSample results from these artifacts.
Inference Hyperparameters  We investigate CFG, noise scale and the number diffusion steps. In every experiment, we fix all values to the default inference values following those described in Section V, and change only a single hyperparameter. The CFG linear extrapolation coefficient increases the speaker similarity, but degrades the automatic quality metric, as seen in Figure 6(a). Next, we scale the noise level added in each diffusion step by scaling the $\beta_{t}\epsilon$ term in Equation 2, and see improvements in similarity but degradation in the UTMOS quality score (Figure 6(b)). We also examine the number of diffusion steps, which improve similarity until reaching 20 diffusion steps, and also degrade UTMOS (Figure 6(c)).